OpenStudy (anonymous):

could you explain why the logarithm of 1 with base b is 0?

OpenStudy (anonymous):

log(base b)1=0, let us raise the power of both sides of the equation like this b^(log(base b)1=b^0 1=1 true

OpenStudy (anonymous):

any question boocoo?

OpenStudy (anonymous):

no, thank you very much! :)

OpenStudy (anonymous):

ok, yw, good luck